OpenStudy (anonymous):

9, 5, 1, -3 Which of the following functions best defines this sequence?

OpenStudy (anonymous):

f(1) = 9, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 9, f(n + 1) = f(n) + 4; for n ≥ 1 f(1) = 9, f(n + 1) = f(n) - 5; for n ≥ 1 f(1) = 9, f(n + 1) = f(n) + 5; for n ≥ 1